OpenStudy (anonymous):

WILL GIVEMEDAL, Which equation can be used to solve the problem? How many 8-gallon cans are needed to hold a total of 120 gallons of water? A. c ÷ 8 = 120 B. 8 ÷ c = 120 C. 8 + c = 120 D. 8c = 120

OpenStudy (danjs):

how many 8 gallon cans does it take to fill a 120 gallon resevoir?

OpenStudy (danjs):

8 gallons times some number 'x' equals 120 gallons

OpenStudy (anonymous):

not sure

OpenStudy (danjs):

In this problem they use c c = # of cans

OpenStudy (danjs):

so 8 times the number of cans 'c' will = 120 total gallons

OpenStudy (danjs):

8*c = 120

OpenStudy (anonymous):

so its c

OpenStudy (danjs):

if you wanted to solve and figure how many cans it would take, divide both sides of that by 8, c = 120/8 = 15

OpenStudy (danjs):

D) 8c = 120

OpenStudy (danjs):

"8 times the number of cans 'c' is 120 gallons"

OpenStudy (anonymous):

d?

OpenStudy (danjs):

yeah it is D 8 times C = 120

OpenStudy (danjs):

Think about it, you want to divide that 120 gallons into 8 gallon parts c = 120/8 or 8c = 120